# A conjecture of Radu and Sellers on congruences modulo powers of 2 for broken 3-diamond partitions

### v1

@@ 4. Radu and Sellers’ conjecture
(4.22) $\pi(\sum_{k,u,r,v}x_{k,u}(m,i)x_{r,v}(n,j)D^{(2s-1)}(l_{k,u},l_{r,v}))$
$\geq$ $\min_{(k,u)\neq(r,v)}(k+r+2s-2+\lambda_{u,v}+\lfloor\frac{2m-k+\gamma_{u,i}}{2}\rfloor+\lfloor\frac{2n-r+\gamma_{v,j}}{2}\rfloor)$
$=$ $\min_{(k,u)\neq(r,v)}(m+n+2s-2+\lambda_{u,v}+\lfloor\frac{k+\gamma_{u,i}}{2}\rfloor+\lfloor\frac{r+\gamma_{v,j}}{2}\rfloor)$
$\geq$ $2s-1+m+n+\lambda_{i,j}.$
By (4.20)–(4.22) we get that inequality (4.17) holds for $\alpha=s$. For $\pi(D^{(2s)}(l_{1,1},l_{3,3}))$ and $\pi(D^{(2s)}(l_{2,1},l_{3,3}))$, based on Lemma 3.3 we derive:
@@
Rewriting $\pi(D^{(2s)}(l_{1,1},l_{3,3}))$ into $\pi(D^{(2s-1)}(4l_{1,0}+l_{2,0}+35l_{1,2}+7l_{2,2},-4l_{1,0}-l_{2,0}-36l_{1,2}-8l_{2,2}))$, we get $\pi(D^{(2s)}(l_{1,1},l_{3,3}))\geq 2s$ by (4.14) and (4.15). Similarly, we obtain $\pi(D^{(2s)}(l_{2,1},l_{3,3}))\geq 2s+1$ by (4.14).
Finally, we assume that (4.14)–(4.19) hold for $\alpha=s$ and prove that (4.14)–(4.16) hold for $\alpha=s+1$. For $u=1$ or $3$, define $y_{k,u}(m,i):=[p_{i}t^{m}]U^{(0)}(p_{u}t^{k})$ be the coefficient of $p_{i}t^{m}$ in $U^{(0)}(p_{u}t^{k})$, namely
@@
With the help of (4.5) and (4.6), we have $\pi(y_{k,u}(m,i))\geq\lfloor\frac{2m-k+\gamma_{u,i}}{2}\rfloor$, where $\gamma_{1,0}=1$, $\gamma_{1,2}=-1$, $\gamma_{3,0}=2$, and $\gamma_{3,2}=2$. Apart from $(k,u,r,v)\in\{(1,1,3,3),(2,1,3,3),(3,3,1,1),(3,3,2,1)\}$, we obtain that
(4.27) $\pi(\sum_{k,u,r,v}y_{k,u}(m,i)y_{r,v}(n,j)D^{(2s)}(l_{k,u},l_{r,v}))$
$\geq$ $\min_{(k,u)\neq(r,v)}(m+n+2s-1+\lambda_{u,v}-\delta(k,u,r,v)+\lfloor\frac{k+\gamma_{u,i}}{2}\rfloor+\lfloor\frac{r+\gamma_{v,j}}{2}\rfloor)$
$\geq$ $2s+m+n+\lambda_{i,j},$
where the last inequality is determined by (4.17) when $\alpha=s$. By (4.18) and (4.19) we have
(4.28) $\pi(y_{1,1}(m,i)y_{3,3}(n,j)D^{(2s)}(l_{1,1},l_{3,3}))\geq 2s+m+n-2,$
(4.29) $\pi(y_{2,1}(m,i)y_{3,3}(n,j)D^{(2s)}(l_{2,1},l_{3,3}))\geq 2s+m+n-2.$
Hence (4.14) holds for $\alpha=s+1$ by (4.26)–(4.29). Next, we consider (4.15) and (4.16). By Lemma 3.3, we get that
@@
Rewriting $\pi(D^{(2s+1)}(l_{1,2},l_{2,0}))$ by (4.30)–(4.33). Extracting terms which are less than $22s+$ 2 by (4.17) and (4.18), we get
@@
$l_{5,3}^{(2s)}=$ $-208l_{1,0}^{(2s-1)}-180l_{2,0}^{(2s-1)}-96l_{3,0}^{(2s-1)}-34l_{4,0}^{(2s-1)}-8l_{5,0}^{(2s-1)}-l_{6,0}^{(2s-1)}-2748l_{1,2}^{(2s-1)}$
$-1944l_{2,2}^{(2s-1)}-926l_{3,2}^{(2s-1)}-306l_{4,2}^{(2s-1)}-68l_{5,2}^{(2s-1)}-8l_{6,2}^{(2s-1)},$
which again derived by Lemma 3.3, we obtain that $\pi(D^{(2s)}(l_{3,3}+l_{1,1},l_{5,3}+2l_{2,1}))\geq 2s+2$ by $\pi(D^{(2s-1)}(l_{1,2},2l_{2,2}+l_{3,2}))\geq 2s$ and $\pi(D^{(2s-1)}(l_{1,2},l_{2,0}))\geq 2s$. Thus, $\pi(D^{(2s+1)}(l_{1,2},l_{2,0}))\geq 2s+2$, which means that (4.15) holds for $\alpha=s+1$. Similarly, rewriting $\pi(D^{(2s+1)}(l_{1,2},2l_{2,2}+l_{3,2}))$ by (4.30)–(4.32), we attain that each term is not less than $22s+$ 2 by (4.17) so that $\pi(D^{(2s+1)}(l_{1,2},2l_{2,2}+l_{3,2}))\geq 2s+2$. Thus, we prove that (4.15) and (4.16) hold for $\alpha=s+1$. ∎
@@
For each $\alpha\geq 1$ there exist a constant $x_{\alpha}$ such that
@@
$\pi(D^{(2\alpha-1)}(l_{m,i},l_{n,j}))\geq 2\alpha-2,$
$\pi(D^{(2\alpha)}(l_{m,i},l_{n,j}))\geq 2\alpha.$
@@
(4.34) $U^{(1)}(U^{(0)}(p_{1}t))\equiv p_{1}t\pmod{2},~{}~{}~{}\quad~{}~{}~{}U^{(0)}(U^{(1)}(y_{0}t))\equiv y_{0}t\pmod{2}.$
Using (4.9) and (4.10), we have $L_{1}\equiv l_{1,2}^{(1)}y_{0}t\equiv y_{0}t\pmod{2}$ and $L_{2}\equiv l_{1,1}^{(2)}p_{1}t\equiv p_{1}t\pmod{2}$, where $l_{1,2}^{(1)}=1$ and $2\nmid l_{1,1}^{(2)}$. Then, by (4.34), we have for $\alpha\geq 1$,
$L_{2\alpha-1}\equiv l_{1,2}^{(2\alpha-1)}y_{0}t\pmod{2},~{}~{}~{}\quad~{}~{}~{}L_{2\alpha}\equiv l_{1,1}^{(2\alpha)}p_{1}t\pmod{2}.$
Note that $2\nmid l_{1,1}^{(2\alpha)}$ and $2\nmid l_{1,2}^{(2\alpha-1)}$. Let $x_{2\alpha-1}$ be a solution of $l_{1,2}^{(2\alpha-1)}\equiv x_{2\alpha-1}l_{1,2}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}$, then for $(n,i)\neq(1,2)$,
@@
Cancelling $l_{1,2}^{(2\alpha-1)}$, we obtain $l_{n,i}^{(2\alpha-1)}\equiv x_{2\alpha-1}l_{n,i}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}$. Similarly, let $x_{2\alpha}$ be a solution of $l_{1,1}^{(2\alpha)}\equiv x_{2\alpha}l_{1,1}^{(2\alpha+2)}\pmod{2^{2\alpha}}$, then for $(n,i)\neq(1,1)$,
@@
$\Delta_{3}(\lambda_{\alpha})\Delta_{3}(2^{\alpha+2}n+\lambda_{\alpha+2})\equiv\Delta_{3}(\lambda_{\alpha+2})\Delta_{3}(2^{\alpha}n+\lambda_{\alpha})\pmod{2^{2\lfloor\frac{\alpha}{2}\rfloor}},$
@@
$+1211\cdot 2^{3}t^{4}+443\cdot 2^{2}t^{3}+91\cdot 2t^{2}+7t)-2^{15}t^{13}-2^{17}t^{12}-111\cdot 2^{11}t^{11}$
$-245\cdot 2^{10}t^{10}-1495\cdot 2^{7}t^{9}-817\cdot 2^{7}t^{8}-5\cdot 2^{13}t^{7}-1387\cdot 2^{3}t^{6}-243\cdot 2^{3}t^{5}-93\cdot 2t^{4}$
$-2^{3}t^{3}.$

### v2

@@ 4. Radu and Sellers’ conjecture
$\pi\!\left(\sum_{k,u,r,v}x_{k,u}(m,i)x_{r,v}(n,j)D^{(2s-1)}(l_{k,u},l_{r,v})\right)$
$\qquad\geq\min_{(k,u)\neq(r,v)}\left(k+r+2s-2+\lambda_{u,v}+\left\lfloor\frac{2m-k+\gamma_{u,i}}{2}\right\rfloor+\left\lfloor\frac{2n-r+\gamma_{v,j}}{2}\right\rfloor\right)$
$\qquad=\min_{(k,u)\neq(r,v)}\left(m+n+2s-2+\lambda_{u,v}+\left\lfloor\frac{k+\gamma_{u,i}}{2}\right\rfloor+\left\lfloor\frac{r+\gamma_{v,j}}{2}\right\rfloor\right)$
(4.22) $\qquad\geq 2s-1+m+n+\lambda_{i,j}.$
By (4.20)–(4) we obtain that inequality (4.17) holds for $\alpha=s$. For $\pi\bigl(D^{(2s)}(l_{1,1},l_{3,3})\bigr)$ and $\pi\bigl(D^{(2s)}(l_{2,1},l_{3,3})\bigr)$, based on Lemma 3.3 we derive:
@@
Rewriting $\pi\bigl(D^{(2s)}(l_{1,1},l_{3,3})\bigr)$ as $\pi\bigl(D^{(2s-1)}(4l_{1,0}+l_{2,0}+35l_{1,2}+7l_{2,2},\;-4l_{1,0}-l_{2,0}-36l_{1,2}-8l_{2,2})\bigr)$, we obtain $\pi\bigl(D^{(2s)}(l_{1,1},l_{3,3})\bigr)\geq 2s$ by (4.14) and (4.15). Similarly, we obtain $\pi\bigl(D^{(2s)}(l_{2,1},l_{3,3})\bigr)\geq 2s+1$ by (4.14).
Finally, we assume that (4.14)–(4.19) hold for $\alpha=s$ and prove that (4.14)–(4.16) hold for $\alpha=s+1$. For $u=1$ or $3$, define $y_{k,u}(m,i):=[p_{i}t^{m}]U^{(0)}(p_{u}t^{k})$ to be the coefficient of $p_{i}t^{m}$ in $U^{(0)}(p_{u}t^{k})$, namely
@@
With the help of (4.5) and (4.6), we have $\pi\bigl(y_{k,u}(m,i)\bigr)\geq\left\lfloor\frac{2m-k+\gamma_{u,i}}{2}\right\rfloor$, where $\gamma_{1,0}=1$, $\gamma_{1,2}=-1$, $\gamma_{3,0}=2$, and $\gamma_{3,2}=2$. Apart from $(k,u,r,v)\in\{(1,1,3,3),(2,1,3,3),(3,3,1,1),(3,3,2,1)\}$, we obtain that
$\pi\!\left(\sum_{k,u,r,v}y_{k,u}(m,i)y_{r,v}(n,j)D^{(2s)}(l_{k,u},l_{r,v})\right)$
$\qquad\geq\min_{(k,u)\neq(r,v)}\left(m+n+2s-1+\lambda_{u,v}-\delta(k,u,r,v)+\left\lfloor\frac{k+\gamma_{u,i}}{2}\right\rfloor+\left\lfloor\frac{r+\gamma_{v,j}}{2}\right\rfloor\right)$
(4.27) $\qquad\geq 2s+m+n+\lambda_{i,j},$
where the last inequality is determined by (4.17) for $\alpha=s$. By (4.18) and (4.19) we have
(4.28) $\pi\!\left(y_{1,1}(m,i)y_{3,3}(n,j)D^{(2s)}(l_{1,1},l_{3,3})\right)\geq 2s+m+n-2,$
(4.29) $\pi\!\left(y_{2,1}(m,i)y_{3,3}(n,j)D^{(2s)}(l_{2,1},l_{3,3})\right)\geq 2s+m+n-2.$
Hence (4.14) holds for $\alpha=s+1$ by (4.26)–(4.29).
Next, we consider (4.15) and (4.16). By Lemma 3.3, we obtain
@@
Rewriting $\pi\bigl(D^{(2s+1)}(l_{1,2},l_{2,0})\bigr)$ using (4.30)–(4.33) and extracting terms that are less than $2s+2$ by (4.17) and (4.18), we obtain
@@
$l_{5,3}^{(2s)}$ $=-208l_{1,0}^{(2s-1)}-180l_{2,0}^{(2s-1)}-96l_{3,0}^{(2s-1)}-34l_{4,0}^{(2s-1)}-8l_{5,0}^{(2s-1)}-l_{6,0}^{(2s-1)}$
$\quad-2748l_{1,2}^{(2s-1)}-1944l_{2,2}^{(2s-1)}-926l_{3,2}^{(2s-1)}-306l_{4,2}^{(2s-1)}-68l_{5,2}^{(2s-1)}-8l_{6,2}^{(2s-1)},$
which is again derived by Lemma 3.3, we obtain that $\pi\bigl(D^{(2s)}(l_{3,3}+l_{1,1},\;l_{5,3}+2l_{2,1})\bigr)\geq 2s+2$ by $\pi\bigl(D^{(2s-1)}(l_{1,2},2l_{2,2}+l_{3,2})\bigr)\geq 2s$ and $\pi\bigl(D^{(2s-1)}(l_{1,2},l_{2,0})\bigr)\geq 2s$. Thus, $\pi\bigl(D^{(2s+1)}(l_{1,2},l_{2,0})\bigr)\geq 2s+2$, which means that (4.15) holds for $\alpha=s+1$. Similarly, rewriting $\pi\bigl(D^{(2s+1)}(l_{1,2},2l_{2,2}+l_{3,2})\bigr)$ using (4.30)–(4), we find that each term is at least $2s+2$ by (4.17), so that $\pi\bigl(D^{(2s+1)}(l_{1,2},2l_{2,2}+l_{3,2})\bigr)\geq 2s+2$. Thus, we prove that (4.15) and (4.16) hold for $\alpha=s+1$. ∎
@@
For each $\alpha\geq 1$ there exists an integral constant $x_{\alpha}$ such that
@@
$\pi\bigl(D^{(2\alpha-1)}(l_{m,i},l_{n,j})\bigr)$ $\geq 2\alpha-2,$
$\pi\bigl(D^{(2\alpha)}(l_{m,i},l_{n,j})\bigr)$ $\geq 2\alpha.$
@@
(4.34) $U^{(1)}(U^{(0)}(p_{1}t))\equiv p_{1}t\pmod{2},\qquad U^{(0)}(U^{(1)}(y_{0}t))\equiv y_{0}t\pmod{2}$.
Using (4.9) and (4.10), we obtain $L_{1}\equiv l_{1,2}^{(1)}y_{0}t\equiv y_{0}t\pmod{2}$ and $L_{2}\equiv l_{1,1}^{(2)}p_{1}t\equiv p_{1}t\pmod{2}$, where $l_{1,2}^{(1)}=1$ and $2\nmid l_{1,1}^{(2)}$.
From (4.34), for $\alpha\geq 1$, we have
$L_{2\alpha-1}\equiv l_{1,2}^{(2\alpha-1)}y_{0}t\equiv y_{0}t\pmod{2},\qquad L_{2\alpha}\equiv l_{1,1}^{(2\alpha)}p_{1}t\equiv p_{1}t\pmod{2}$.
This implies that $2\nmid l_{1,1}^{(2\alpha)}$ and $2\nmid l_{1,2}^{(2\alpha-1)}$. Let $x_{2\alpha-1}$ be a solution of $l_{1,2}^{(2\alpha-1)}\equiv x_{2\alpha-1}l_{1,2}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}$. Then for $(n,i)\neq(1,2)$,
@@
Cancelling $l_{1,2}^{(2\alpha-1)}$, we obtain $l_{n,i}^{(2\alpha-1)}\equiv x_{2\alpha-1}l_{n,i}^{(2\alpha+1)}\pmod{2^{2\alpha-2}}$. Similarly, let $x_{2\alpha}$ be a solution of $l_{1,1}^{(2\alpha)}\equiv x_{2\alpha}l_{1,1}^{(2\alpha+2)}\pmod{2^{2\alpha}}$. Then for $(n,i)\neq(1,1)$,
@@
$\Delta_{3}(\lambda_{\alpha})\,\Delta_{3}(2^{\alpha+2}n+\lambda_{\alpha+2})\equiv\Delta_{3}(\lambda_{\alpha+2})\,\Delta_{3}(2^{\alpha}n+\lambda_{\alpha})\pmod{2^{2\lfloor\alpha/2\rfloor}}$,
@@
$\lambda_{\alpha}=\left\{\begin{array}[]{ll}\dfrac{2^{\alpha+1}+1}{3},&\text{if }\alpha\text{ is even},\\[6.0pt] \dfrac{2^{\alpha}+1}{3},&\text{if }\alpha\text{ is odd}.\end{array}\right$.
@@
$+1211\cdot 2^{3}t^{4}+443\cdot 2^{2}t^{3}+91\cdot 2t^{2}+7t)-2^{15}t^{13}-2^{17}t^{12}-111\cdot 2^{11}t^{11}-245\cdot 2^{10}t^{10}$
$-1495\cdot 2^{7}t^{9}-817\cdot 2^{7}t^{8}-5\cdot 2^{13}t^{7}-1387\cdot 2^{3}t^{6}-243\cdot 2^{3}t^{5}-93\cdot 2t^{4}-2^{3}t^{3}.$
@@ Acknowledgements
The second author was supported by the National Natural Science Foundation of China (Grant No. 12401438).